The smallest integer in a set of three consecutive integers is x, and the largest integer in this set is z. If the sum of the set is equal to 27, then what is the value of xz?
Answer by mananth(16946)

Let the smallest integer be x.
Since the three integers are consequtive
the next two integers will be x+1 , and x+2.
The sum of all the three consecutive integers = 27. ( given in the problem)
x + x+1 + x+2 = 27
3x + 3= 27
3x= 27-3
3x=24
x=24/3 =8
the three consequitve integers are 8, 8+1, 8+2
ie: 8, 9 , 10.
Z= largest integer = 10 given in the problem.
x=8 the smallest integer.
x * z = 8* 10 = 80